# Incorporating Flexible Image Conditioning into Text-to-Video Diffusion Models without Training

### v1

@@ Appendix C Implementation Details
Table 4: Prompt for each selected action class in UCF-101. More details of the test set is elaborated in Sec. C.1.
@@ C.1 Establishment of Test Set
Ni et al. [37] select 10 classes in UCF-101 which are not diverse enough for a reliable evaluation. The textual prompts used in their work are also oversimplified. In this work, we further extend the test set to 42 classes with manually annotated textual prompt for each action. For each class, there are videos of 25 subjects performing the same action. For diversity, we randomly select one video of each subject in each action class, totaling 1050 test data samples. The prompts are listed in Tab. 4.
@@ C.2 More Details of Our Method
In the proposed method, the model also conditions on textual prompts. We adopt classifier-free guidance in the inference with guidance scale as 9.0. We use a deterministic DDIM sampling strategy ($\sigma_{t}=0$ in main paper Eq. 2) that can accelerate generation by reducing denoising steps [45]. We generate videos with $M=16$ frames in all experiments for a fair comparison with previous methods. Note that our method can naturally extend to videos with more frames. Our method is implemented on a single NVIDIA H100 GPU.
@@ C.3 Implementation of Previous Methods
We implement all baseline models using their publicly released codebases following the default hyperparameter settings. To adapt TI2V-Zero [37] to rewinding and outpainting tasks, we reverse the order of frame generation for rewinding, and generate the frames before and after the condition image in two directions separately for outpainting. For TRF [7], we use the setting of $gym+$ in their code.
Appendix D Limitation and Future Work
Though our approach achieves great performance in many experimental settings compared with previous methods, we still observe some limitations as illustrated in Fig. 11. First of all, FlexTI2V is still weak in understanding the transition of camera viewpoints between two condition images. The output video looks like a transition of two slides if there is a big viewpoint change. Second, our method is affected by some flaws in the pretrained foundation models (e.g., watermark) because of the training-free nature. The upperbound of our method is constrained by the T2V model we used. Third, FlexTI2V is able to make the condition images as exact video frames at expected positions. Our method can not synthesize videos following only the subjects or background in the condition images (e.g., personalized video generation).
@@ Appendix D Limitation and Future Work
Appendix E Potential Societal Impacts of Our Research
Positive Societal Impacts
The proposed method has the potential to enhance creative expression and communication in several ways:
  • Creative Storytelling: Artists, filmmakers, and content creators could use this technology to seamlessly integrate specific images in storybooks into dynamically generated videos, reducing the manual effort in video creation.
  • Personalized Video Generation: Individuals could create personalized videos for entertainment, education, or social sharing, for instance, by inserting photos of friends and family into animated stories.
  • Skill Learning: Educators could develop more engaging and customized learning materials by generating videos that include specific diagrams, historical images, or relevant visual cues at precise moments within an explanation.
Negative Societal Impacts
The ability to precisely control the appearance of specific images within generated video content also carries some risks.
  • Fake News: The most direct and alarming negative application is the creation of highly convincing "deepfake" videos. Our method could be exploited to insert images of public figures or private individuals into fabricated scenarios, making them appear to say or do things they never did.
  • Bias Amplification: Our method relies on pretrained T2V foundation models. These models can inherit and amplify societal biases present in their vast training datasets (e.g., associating certain demographics with specific roles, actions, or environments). Since our method does not involve finetuning, the generated video content around the conditioned images will be influenced by these underlying biases. This could lead to the creation of videos that perpetuate harmful stereotypes based on the content of the input images or the accompanying text prompts.
Potential Solutions
  • It is crucial to invest in and advance research into robust methods for detecting AI-generated or manipulated videos. This includes developing tools that can identify synthetic artifacts and exploring resilient watermarking or digital provenance techniques to help distinguish authentic media from generated content.
  • We should encourage continued auditing of foundation models for biases and the development of techniques to mitigate these biases.
Appendix F Code and Data Release

### v2

@@ C Implementation Details
Table E: Prompt for each selected action class in UCF-101. More details of the test set is elaborated in Sec. C.1.
@@ C.1 Establishment of Test Set
ni2024ti2v select 10 classes in UCF-101 which are not diverse enough for a reliable evaluation. The textual prompts used in their work are also oversimplified. In this work, we further extend the test set to 42 classes with manually annotated textual prompt for each action. For each class, there are videos of 25 subjects performing the same action. For diversity, we randomly select one video of each subject in each action class, totaling 1050 test data samples. The prompts are listed in Tab. E.
@@ C.2 More Details of Our Method
In the proposed method, the model also conditions on textual prompts. We adopt classifier-free guidance in the inference with guidance scale as 9.0. We use a deterministic DDIM sampling strategy ($\sigma_{t}=0$ in main paper Eq. 2) that can accelerate generation by reducing denoising steps [song2020denoising]. We generate videos with $M=16$ frames in all experiments for a fair comparison with previous methods. Note that our method can naturally extend to videos with more frames. Our method is implemented on a single NVIDIA H100 GPU.
@@ C.3 Implementation of Previous Methods
We implement all baseline models using their publicly released codebases following the default hyperparameter settings. To adapt TI2V-Zero [ni2024ti2v] to rewinding and outpainting tasks, we reverse the order of frame generation for rewinding, and generate the frames before and after the condition image in two directions separately for outpainting. For TRF [feng2024explorative], we use the setting of $\verb+gym$+ in their code.
D Limitation and Future Work
Though our approach achieves great performance in many experimental settings compared with previous methods, we still observe some limitations as illustrated in Fig. E. First of all, FlexTI2V is still weak in understanding the transition of camera viewpoints between two condition images. The output video looks like a transition of two slides if there is a big viewpoint change. Second, our method is affected by some flaws in the pretrained foundation models (e.g., watermark) because of the training-free nature. The upperbound of our method is constrained by the T2V model we used. Third, FlexTI2V is able to make the condition images as exact video frames at expected positions. Our method can not synthesize videos following only the subjects or background in the condition images (e.g., personalized video generation).
@@ D Limitation and Future Work
E Code and Data Release